OpenStudy (anonymous):

solve f(x)=2 by the function f(x)=x^3-3x^2+3x

OpenStudy (candyme):

This problem is asking you what x is when f(x) is 2. Start by setting your equation equal to 2.

OpenStudy (anonymous):

solve x^3-3x^2+3x-2=0

OpenStudy (anonymous):

solved it.so the formula was x^3-3x^2+3x x=3x^2-6x=3x f(x)=6x-6+3 f'(x)=6x-6=0+3 0+3=3 thus,f'(x)=3

OpenStudy (anonymous):

there may have been a bit of info i left out.it was a table with all the x,f(x) and f'(x) values.

OpenStudy (anonymous):

i had to solve f'(x) when f(x)=2 while using the function mentioned.

OpenStudy (anonymous):

hey,guys,could you help me out with understanding the question?

OpenStudy (ash2326):

Hey @nck93 we are given f(x)=2 and f(x) is defined as x^3-3x^2+3x so \[f(x)=x^3-3x^2+3x\] and for a particular value of x f(x) is 2

OpenStudy (anonymous):

yeah,i get that.i understand solving x and f(x), but not f'(x).like in the attachment the value for f'(x) is not given,but the value for for f(x) is 2 the same on the left side,but with 0,how do i solve that?

OpenStudy (ash2326):

Wait I'm going through attachement

OpenStudy (ash2326):

x=0 we have \[ f(x)=x^3-4x^2+3x\] put x=0 we get \[f(0)=0\] that's why table mentions f(x)=0 for x=0 now differentiate with respect to x \[f'(x)=3x^2-8x+3\] put x=0 to find f'(x) at x=0

OpenStudy (anonymous):

ah,now i get it.thanks alot =)

OpenStudy (ash2326):

welcome:)

OpenStudy (anonymous):

im using 3x^2-8x+3 as 3(0)^2-8(0)+3 is that right?

OpenStudy (ash2326):

Yeah:)

OpenStudy (anonymous):

sorry 3(0) i think you noted the question down wrong. its +3x in the end,not 3.

OpenStudy (anonymous):

oh wait,my mistake.that after differentiation..

OpenStudy (anonymous):

its f(x)= x^3-3X^2+3x you wrote x^3-4x^2+3x

OpenStudy (ash2326):

Sorry \[f'(x)=3x^2-6x+3\] now put x=0 to find f'(x) at x=0

OpenStudy (anonymous):

Np,thanks.

OpenStudy (ash2326):

welcome:)